OpenStudy (anonymous):

A^2+X^2= sqrtx^2+4. Solve for A WILL MEDAL AND FAN

OpenStudy (solomonzelman):

you mean that there is a square root of left side, or the square root (x^2+4) is really squared.

OpenStudy (anonymous):

\[A^2+x^2=\sqrt{x^2+4}\]

OpenStudy (solomonzelman):

A^2 + X^2 = X^2+4 A^2=4 A=2 that would be the adjacent side (Assuming that this is the contiuation of the previous problem)

OpenStudy (anonymous):

yes sir.

OpenStudy (solomonzelman):

And so the adjacent side is 2, and hypotenuse is \(\sqrt{x^2+4}\), then sec would be equivalent to?

OpenStudy (anonymous):

how does A = 2?

OpenStudy (anonymous):

sec is equal to cos

OpenStudy (solomonzelman):

No, sec is the reciprocal of cos

OpenStudy (anonymous):

oops i was reading something else

OpenStudy (solomonzelman):

and A=2, is what we had from: A^2 + X^2 = X^2+4

OpenStudy (solomonzelman):

secant = \(\rm Hypotenuse/Adjacent=\sqrt{x^2+4}/2\)

OpenStudy (anonymous):

we are solving for "a" though. I thought we are suppose to single it out.

OpenStudy (solomonzelman):

and same thing you would get from: \(\large\color{black}{ \displaystyle \frac{1}{\sqrt{\dfrac{x^2+4}{x^2+4}-\dfrac{x^2}{x^2+4}}}=}\) \(\large\color{black}{ \displaystyle \frac{1}{\sqrt{\dfrac{4}{x^2+4}}}=}\) \(\large\color{black}{ \displaystyle \frac{1}{\dfrac{2}{\sqrt{x^2+4}}}=}\) \(\large\color{black}{ \displaystyle \frac{\sqrt{x^2+4}}{2}.}\)

OpenStudy (anonymous):

Can you do it where you would single out the "A"

OpenStudy (solomonzelman):

I have solved for A, haven't I showed how?

OpenStudy (solomonzelman):

you had that: \(\large\color{black}{ \displaystyle {\rm A}^2+x^2=\left(\sqrt{x^2+4}\right)^2 }\)

OpenStudy (anonymous):

i know you solved for A. but i dont understand how "A" becomes 2

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm A}^2+x^2=\left(\sqrt{x^2+4}\right)^2 }\) \(\large\color{black}{ \displaystyle {\rm A}^2+x^2=|x^2+4| }\) assuming that x is real, \(\large\color{black}{ \displaystyle {\rm A}^2+x^2=x^2+4 }\) \(\large\color{black}{ \displaystyle {\rm A}^2=4\quad \Longleftarrow \quad {\rm A}=2 }\)

OpenStudy (anonymous):

omg im so sorry i over complicated this lol. i got it now.

OpenStudy (solomonzelman):

it's fine, \(....\) (And same thing you will get for the secant, if you do what I initally offered.)

OpenStudy (anonymous):

ok thanks a ton